Integral 0 to pi e^cosx cos (sinx) dx

Grade:11

1 Answers

Aravind Bommera
11 years ago

Hi,Shreya

 

Let u = sin(x). Then du = cos(x) dx. So you can now antidifferentiate e^u du. This
is e^u + C = e^sin(x) + C.

 

Then substitute your range 0 to pi.

e^sin (pi)-e^sin(0)

=0-0

=0